# Learning Energy-Based Models by Cooperative Diffusion Recovery Likelihood

### v3

@@ 3.2 Amortizing MCMC sampling with initializer models
To address this issue, we propose to learn an initializer model jointly with the EBM at each noise level, which maps ${\mathbf{x}}_{t+1}$ closer to the manifold of ${\mathbf{y}}_{t}$. Our work is inspired by the CoopNets work Xie et al. (2018a; 2021b; 2022b), which shows that jointly training a top-down generator via MCMC teaching will help the training of a single EBM model. We take this idea and generalize it to the recovery-likelihood model. More discussions are included in Appendix A. Specifically, the initializer model at noise level $t$ is defined as
@@
It serves as a coarse approximation to $p_{\theta}({\mathbf{y}}_{t}|{\mathbf{x}}_{t+1})$, as the former is a single-mode Gaussian distribution while the latter can be multi-modal. A more general formulation would be to involve latent variables ${\mathbf{z}}_{t}$ following a certain simple prior $p({\mathbf{z}}_{t})$ into ${\bm{g}}_{\phi}$. Then $q_{\phi}({\mathbf{y}}_{t},t|{\mathbf{x}}_{t+1})=\mathbb{E}_{p({\mathbf{z}}_{t})}\left[q_{\phi}({\mathbf{y}}_{t},{\mathbf{z}}_{t},t|{\mathbf{x}}_{t+1})\right]$ can be non-Gaussian (Xiao et al., 2022). However, we empirically find that the simple initializer in Equation 6 works well. Compared with the more general formulation, the simple initializer avoids the inference of ${\mathbf{z}}_{t}$ which may again require MCMC sampling, and leads to more stable training. Different from (Xiao et al., 2022), samples from the initializer just serves as the starting points and are refined by sampling from the EBM, instead of being treated as the final samples. We follow (Ho et al., 2020) to set $\tilde{\sigma}_{t}=\sqrt{\frac{1-\bar{\alpha}_{t}^{2}}{1-\bar{\alpha}_{t+1}^{2}}}\sigma_{t}$. If we treat the sequence of initializers as the reverse process, such choice of $\tilde{\sigma}_{t}$ corresponds to the lower bound of the standard deviation given by $p_{\rm data}$ being a delta function (Sohl-Dickstein et al., 2015).
@@ 3.5 Conditional generation and classifier-free guidance
(Ho & Salimans, 2022) proposed classifier-free guidance that greatly improves the sample quality of conditional diffusion models, and trades-off between sample quality and sample diversity by adjusting the guidance strength. Given the close connection between EBMs and diffusion models, we show that it is possible to apply classifier-free guidance in CDRL as well. Specifically, suppose $c$ is the context (e.g., a label or a text description). At each noise level we jointly estimate an unconditional EBM $p_{\theta}({\mathbf{y}}_{t})\propto\exp(f_{\theta}({\mathbf{y}}_{t};t))$ and a conditional EBM $p_{\theta}({\mathbf{y}}_{t}|c)\propto\exp(f_{\theta}({\mathbf{y}}_{t};c,t))$. By Bayes rule:
$p_{\theta}(c|{\mathbf{y}}_{t})=\frac{p_{\theta}(c,{\mathbf{y}}_{t})}{p_{\theta}({\mathbf{y}}_{t})}=\frac{p_{\theta}({\mathbf{y}}_{t}|c)p(c)}{p_{\theta}({\mathbf{y}}_{t})}.$ (10)
With classifier-free guidance, we assume that the log-density of ${\mathbf{y}}_{t}$ is scaled to
$\log\tilde{p}_{\theta}({\mathbf{y}}_{t}|c)=\log\left[p_{\theta}({\mathbf{y}}_{t}|c)p_{\theta}(c|{\mathbf{y}}_{t})^{w}\right]+{\rm const.}=(w+1)f_{\theta}({\mathbf{y}}_{t};c,t)-wf_{\theta}({\mathbf{y}}_{t};t)+{\rm const.},$ (11)
where $w$ controls the guidance strength. Similarly, for the initializer model, we jointly estimate an unconditional model $q_{\phi}({\mathbf{y}}_{t}|{\mathbf{x}}_{t+1})\sim\mathcal{N}({\bm{g}}_{\phi}({\mathbf{x}}_{t+1};t),\tilde{\sigma}_{t}^{2}{\bm{I}})$ and a conditional model $q_{\phi}({\mathbf{y}}_{t}|c,{\mathbf{x}}_{t+1})\sim\mathcal{N}({\bm{g}}_{\phi}({\mathbf{x}}_{t+1};c,t),\tilde{\sigma}_{t}^{2}{\bm{I}})$. Since both models follow Gaussian distributions, the scaled conditional distribution with classifier-free guidance is still a Gaussian (Dhariwal & Nichol, 2021):
$\tilde{q}_{\phi}({\mathbf{y}}_{t}|c,{\mathbf{x}}_{t+1})\propto q_{\phi}({\mathbf{y}}_{t}|c,{\mathbf{x}}_{t+1})q_{\phi}(c|{\mathbf{y}}_{t},{\mathbf{x}}_{t+1})^{w}\sim\mathcal{N}\left((w+1){\bm{g}}_{\phi}({\mathbf{x}}_{t+1};c,t)-w{\bm{g}}_{\phi}({\mathbf{x}}_{t+1};t),\tilde{\sigma}_{t}^{2}{\bm{I}}\right).$ (12)
@@ 3.6 Compositionality in energy-based model
One attractive property of EBMs is compositionality: one can combine multiple EBMs conditioned on individual concepts, and re-normalize it to create a new distribution conditioned on the intersection of those concepts. Specifically, given two EBMs $p_{\theta}({\mathbf{x}}|c_{1})\propto\exp(f_{\theta}({\mathbf{x}};c_{1}))$ and $p_{\theta}({\mathbf{x}}|c_{2})\propto\exp(f_{\theta}({\mathbf{x}};c_{2}))$ that are conditional on two seperate concepts, (Du et al., 2020; Lee et al., 2023) constructs a new EBM conditional on both concepts as $p_{\theta}({\mathbf{x}}|c_{1},c_{2})\propto\exp(f_{\theta}({\mathbf{x}};c_{1})+f_{\theta}({\mathbf{x}};c_{2}))$ based on the production of expert (Hinton, 2002). Here we show that a negative energy term is missing in that formulation, which can be easily added back with classifier-free guidance and leads to better compositional generation empirically. Specifically, suppose the two concepts $c_{1}$ and $c_{2}$ are conditionally independent given the observed data ${\mathbf{x}}$. Then we have
@@
$=\log p_{\theta}({\mathbf{x}}|c_{1})+\log p_{\theta}({\mathbf{x}}|c_{2})-\log p_{\theta}({\mathbf{x}})+{\rm const.}$
The composition can be generalized to include arbitrary number of concepts. Suppose we have $M$ conditionally independent concepts $c_{i},i=1,...,M$, then
@@
We can combine the compositional log-density (Equation 13) with classifier-free guidance (Equation 11) to further improve the alignment of generated samples with given concepts. The scaled log-density function is given by
$\log\left[p({\mathbf{x}}|c_{i},i=1,...,M)p(c_{i},i=1,...,M|{\mathbf{x}})^{w}\right]$
$=(w+1)\sum_{i=1}^{M}\log p_{\theta}({\mathbf{x}}|c_{i})-(Mw+M-1)\log p({\mathbf{x}})+{\rm const}.$ (14)
@@ 4.3 Conditional synthesis with classifier-free guidance
Models Number of noise level $\times$ Number of MCMC steps FID normal-$\downarrow$
@@ B.4 Illustration of the CDRL Framework
(a) CDRL training process. In the training phase, we start by selecting a pair of images at noise levels $t$ and $t-1$. The image at noise level $t$ is then fed into the initializer to generate an initial proposal. Subsequently, this initial proposal undergoes refinement through the MCMC process guided by the underlying energy function. The refined sample obtained from this process is utilized to update both the energy function and the initializer.
(b) CDRL Sampling process. The sampling phase starts from Gaussian noise. Starting from the highest noise level, an initial proposal is generated by the initializer that corresponds to that noise level. Subsequently, the samples undergo refinement through MCMC sampling. This denoising process is iteratively repeated to push the noisy image towards lower noise levels until the lowest noise level is reached.
@@ D.4 Analyzing the Effects of the Initializer and the EBM
To gain deeper insights into the roles of the initializer and the EBM in the CDRL in image generation, we conduct two additional experiments using a pretrained CDRL model on the ImageNet Dataset ($32\times 32$). We evaluate two generation options: (a) images generated using only the initializer’s proposal, without the EBM’s Langevin Dynamics at each noise level, and (b) images generated with the full CDRL model, which includes the initializer’s proposal and 15-step Langevin updates at each noise level. As shown in Figure 8(a) and 8(b), the initializer captures the rough outline of the object, while the Langevin updates by the EBM improve the details of the object. Furthermore, in Figure 8(c), we display samples generated by fixing the initial noise image and sample noise of each initializer proposal step. The outcomes demonstrate that images generated with the same initialization noises share basic elements but differ in details, highlighting the impact of both the initializer and the Langevin sampling. The initializer provides a starting point, while the Langevin sampling process enriches details.

### v4

@@ 3.2 Amortizing MCMC sampling with initializer models
To address this issue, we propose to learn an initializer model jointly with the EBM at each noise level, which maps ${\mathbf{x}}_{t+1}$ closer to the manifold of ${\mathbf{y}}_{t}$. Our work is inspired by the CoopNets (Xie et al., 2018a; 2021b; 2022b), which shows that jointly training a top-down generator via MCMC teaching will help the training of a single EBM model. We take this idea and generalize it to the recovery-likelihood model. More discussions are included in Appendix A. Specifically, the initializer model at noise level $t$ is defined as
@@
It serves as a coarse approximation to $p_{\theta}({\mathbf{y}}_{t}|{\mathbf{x}}_{t+1})$, as the former is a single-mode Gaussian distribution while the latter can be multi-modal. A more general formulation would be to involve latent variables ${\mathbf{z}}_{t}$ following a certain simple prior $p({\mathbf{z}}_{t})$ into ${\bm{g}}_{\phi}$. Then $q_{\phi}({\mathbf{y}}_{t},t|{\mathbf{x}}_{t+1})=\mathbb{E}_{p({\mathbf{z}}_{t})}\left[q_{\phi}({\mathbf{y}}_{t},{\mathbf{z}}_{t},t|{\mathbf{x}}_{t+1})\right]$ can be non-Gaussian (Xiao et al., 2022). However, we empirically find that the simple initializer in Equation 6 works well. Compared with the more general formulation, the simple initializer avoids the inference of ${\mathbf{z}}_{t}$ which may again require MCMC sampling, and leads to more stable training. Different from Xiao et al. (2022), samples from the initializer just serves as the starting points and are refined by sampling from the EBM, instead of being treated as the final samples. We follow (Ho et al., 2020) to set $\tilde{\sigma}_{t}=\sqrt{\frac{1-\bar{\alpha}_{t}^{2}}{1-\bar{\alpha}_{t+1}^{2}}}\sigma_{t}$. If we treat the sequence of initializers as the reverse process, such choice of $\tilde{\sigma}_{t}$ corresponds to the lower bound of the standard deviation given by $p_{\rm data}$ being a delta function (Sohl-Dickstein et al., 2015).
@@ 3.5 Conditional generation and classifier-free guidance
Ho & Salimans (2022) introduced classifier-free guidance, which greatly enhances the sample quality of conditional diffusion models and balances between sample quality and diversity by adjusting the strength of guidance. Given the close connection between EBMs and diffusion models, we show that it is possible to apply classifier-free guidance in our CDRL as well. Specifically, suppose $c$ is the context (e.g., a label or a text description). At each noise level we jointly estimate an unconditional EBM $p_{\theta}({\mathbf{y}}_{t})\propto\exp(f_{\theta}({\mathbf{y}}_{t};t))$ and a conditional EBM $p_{\theta}({\mathbf{y}}_{t}|c)\propto\exp(f_{\theta}({\mathbf{y}}_{t};c,t))$. By following the classifier-free guidance (Ho & Salimans, 2022), we can perform conditional sampling using the following linear combination of the conditional and unconditional gradients:
$\nabla_{{\mathbf{y}}}\log\tilde{p}_{\theta}({\mathbf{y}}_{t}|c)$ $=(w+1)\nabla_{{\mathbf{y}}}\log p_{\theta}({\mathbf{y}}_{t};c,t)-w\nabla_{{\mathbf{y}}}\log p_{\theta}({\mathbf{y}}_{t};t)$ (10)
$=(w+1)\nabla_{{\mathbf{y}}}f_{\theta}({\mathbf{y}}_{t};c,t)-w\nabla_{{\mathbf{y}}}f_{\theta}({\mathbf{y}}_{t};t),$ (11)
where $w$ controls the guidance strength. We adopt a single neural network to parameterize both conditional and unconditional models, where for the unconditional model we can simply input a null token $\varnothing$ for the class condition $c$ when outputing the negative energy, i.e. $f_{\theta}({\mathbf{y}}_{t};t)=f_{\theta}({\mathbf{y}}_{t};c=\varnothing,t)$. Similarly, for the initializer model, we jointly estimate an unconditional model $q_{\phi}({\mathbf{y}}_{t}|{\mathbf{x}}_{t+1})\sim\mathcal{N}({\bm{g}}_{\phi}({\mathbf{x}}_{t+1};t),\tilde{\sigma}_{t}^{2}{\bm{I}})$ and a conditional model $q_{\phi}({\mathbf{y}}_{t}|c,{\mathbf{x}}_{t+1})\sim\mathcal{N}({\bm{g}}_{\phi}({\mathbf{x}}_{t+1};c,t),\tilde{\sigma}_{t}^{2}{\bm{I}})$. We parameterize both models in a single neural network. Since both models follow Gaussian distributions, the scaled conditional distribution with classifier-free guidance is still a Gaussian distribution (Dhariwal & Nichol, 2021):
$\tilde{q}_{\phi}({\mathbf{y}}_{t}|c,{\mathbf{x}}_{t+1})=\mathcal{N}\left((w+1){\bm{g}}_{\phi}({\mathbf{x}}_{t+1};c,t)-w{\bm{g}}_{\phi}({\mathbf{x}}_{t+1};t),\tilde{\sigma}_{t}^{2}{\bm{I}}\right).$ (12)
@@ 3.6 Compositionality in energy-based model
One attractive property of EBMs is compositionality: one can combine multiple EBMs conditioned on individual concepts, and re-normalize it to create a new distribution conditioned on the intersection of those concepts. Specifically, considering two EBMs $p_{\theta}({\mathbf{x}}|c_{1})\propto\exp(f_{\theta}({\mathbf{x}};c_{1}))$ and $p_{\theta}({\mathbf{x}}|c_{2})\propto\exp(f_{\theta}({\mathbf{x}};c_{2}))$ that are conditioned on two separate concepts $c_{1}$ and $c_{2}$ respectively, Du et al. (2020); Lee et al. (2023) construct a new EBM conditioned on both concepts as $p_{\theta}({\mathbf{x}}|c_{1},c_{2})\propto\exp(f_{\theta}({\mathbf{x}};c_{1})+f_{\theta}({\mathbf{x}};c_{2}))$ based on the production of expert (Hinton, 2002). Specifically, suppose the two concepts $c_{1}$ and $c_{2}$ are conditionally independent given the observed example ${\mathbf{x}}$. Then we have
@@
$=\log p_{\theta}({\mathbf{x}}|c_{1})+\log p_{\theta}({\mathbf{x}}|c_{2})-\log p_{\theta}({\mathbf{x}})+{\rm const.},$
where const. is a constant term independent of ${\mathbf{x}}$. The composition can be generalized to include an arbitrary number of concepts. Suppose we have $M$ conditionally independent concepts, then
@@
We can combine the compositional log-density (Equation 13) with classifier-free guidance (Equation 11) to further improve the alignment of generated samples with given concepts. The sampling gradient of the scaled log-density function is given by
$\nabla_{{\mathbf{x}}}\log\tilde{p}_{\theta}({\mathbf{x}}|c_{i},i=1,...,M)$ $=\nabla_{{\mathbf{x}}}\left[(w+1)\sum_{i=1}^{M}\log p_{\theta}({\mathbf{x}}|c_{i})-(Mw+M-1)\log p_{\theta}({\mathbf{x}})+{\rm const}\right]$
$=(w+1)\sum_{i=1}^{M}\nabla_{{\mathbf{x}}}f_{\theta}({\mathbf{x}}|c_{i})-(Mw+M-1)\nabla_{{\mathbf{x}}}f_{\theta}({\mathbf{x}}).$ (14)
@@ 4.3 Conditional synthesis with classifier-free guidance
Models Number of noise level $\times$ Number of MCMC steps FID $\downarrow$
@@ B.4 Illustration of the CDRL Framework
(a) CDRL training process. In the training phase, we start by selecting a pair of images $\{{\mathbf{y}}_{t},{\mathbf{x}}_{t+1}\}$ at noise levels $t$ and $t+1$ respectively. The image ${\mathbf{x}}_{t}$ at noise level $t$ is then fed into the initializer to generate an initial proposal $\hat{{\mathbf{y}}}_{t}$. Subsequently, this initial proposal undergoes refinement through the MCMC process guided by the underlying energy function. The refined sample $\tilde{{\mathbf{y}}}_{t}$ obtained from this process is utilized to update both the energy function and the initializer.
(b) CDRL Sampling process. The sampling phase starts from Gaussian noise. Starting from the highest noise level, an initial proposal is generated by the initializer that corresponds to that noise level. Subsequently, the samples undergo refinement through MCMC sampling. This denoising process is iteratively repeated to push the noisy image towards lower noise levels until the lowest noise level $t=0$ is reached.
@@ D.4 Analyzing the Effects of the Initializer and the EBM
To gain deeper insights into the roles of the initializer and the EBM in the CDRL in image generation, we conduct two additional experiments using a pretrained CDRL model on the ImageNet Dataset ($32\times 32$). We evaluate two generation options: (a) images generated using only the initializer’s proposal, without the EBM’s Langevin Dynamics at each noise level, and (b) images generated with the full CDRL model, which includes the initializer’s proposal and 15-step Langevin updates at each noise level. As shown in Figure 9(a) and 9(b), the initializer captures the rough outline of the object, while the Langevin updates by the EBM improve the details of the object. Furthermore, in Figure 9(c), we display samples generated by fixing the initial noise image and sample noise of each initializer proposal step. The outcomes demonstrate that images generated with the same initialization noises share basic elements but differ in details, highlighting the impact of both the initializer and the Langevin sampling. The initializer provides a starting point, while the Langevin sampling process enriches details.